# DevFD: Developmental Face Forgery Detection by Learning Shared and Orthogonal LoRA Subspaces

### v2

@@ 4.1 Experimental Setup
Datasets. To facilitate comparisons, we select multiple datasets, employing different forgery methods to manipulate and generate faces from distinct domains. FaceForensics++ (FF++) Rossler et al. (2019) contains 1K real face videos. These faces are manipulated by four forgery methods, resulting in a total of 4K fake videos. The Deepfake Detection (DFD) Dufour and Gully (2019) dataset comprises over a hundred real face samples and more than 1K forged face samples. For the Deepfake Detection Challenge (DFDC) Dolhansky et al. (2020) dataset, we utilized the Preview version (DFDC-P), which includes 5K videos and two forgery methods. Celeb-DF v2 (CDF2) Li et al. (2020b) comprises 590 real and 5,639 fake videos. DF40 Yan et al. (2024) is a large-scale dataset comprising 40 types of forgery methods and over 100K fake videos. We select [FF++, DFDC-P, DFD, CDF2] as our dataset-incremental testing protocol. From DF40, we choose three forgery methods corresponding to different forgery types: Face-Swapping (FS): BlendFace, Face-Reenactment (FR): MCNet, and Entire Face Synthesis (EFS): StyleGAN3. The datasets of selected forgery methods and a base dataset with hybrid forgery types FF++ jointly compose our forgery-type-incremental testing protocol: [Hybrid, FR, FS, EFS].
Implementation Details and Metrics. We utilize a pre-trained ViT model as the initial backbone. The pre-training data is sourced from real faces in the FF++ dataset. We conduct self-supervised pre-training through a data augmentation approach Shiohara and Yamasaki (2022). Regarding data utilization, we strictly isolate data across tasks, including real and fake faces. Similar to DFIL Pan et al. (2023), we randomly sample 100 videos from the current dataset and train for 20 epochs in each task, maintaining a balance between real and fake samples. The hyperparameter settings are adjusted according to the number of training epochs, and detailed information is provided in Appendix C.2. We employed the Adam optimizer with parameters set to $\beta_{1}=0.9,\beta_{2}=0.999$. The learning rate is set to 1e41e-4, and the batch size is set to 128. The rank (subspace dimension) for LoRA was determined using a grid optimization algorithm similar to InfLoRA Liang and Li (2024). We utilize the Accuracy (Acc) and Area Under the Curve (AUC) to evaluate the overall performance of the model and calculate the average forgetting of these two metrics. Average Forgetting (AF) indicates the model’s ability to retain information from previously learned tasks:
$AF=\frac{1}{T}\sum_{i}^{T}\left(Score_{i}^{\text{first }}-Score_{i}^{\text{last }}\right),$ (15)
where $Score$ denotes the evaluation metric (Acc or AUC).
@@ 4.2 Experimental Results
Experiments on Dataset Incremental Protocol. To comprehensively evaluate the performance of the proposed method, we conduct comparative experiments with general continual learning approaches (e.g., LwF Li and Hoiem (2017), ER Chaudhry et al. (2019), and SI Zenke et al. (2017)) and specialized continual face forgery detection methods (e.g., CoReD Kim et al. (2021), DFIL Pan et al. (2023), and DMP Tian et al. (2024)). We ensure that the improvements stem from the new strategy and maintain the fairness of the experiments, as detailed in Appendix C.1. Under the dataset incremental protocol, the model is sequentially trained on task sequences constructed from [FF++, DFDC-P, DFD, CDF2]. After completing each task’s training phase, we evaluate model accuracy on both the current and previously learned tasks. Quantitative results are organized in a lower triangular accuracy matrix, accompanied by two key metrics: average accuracy (Avg) and average forgetting (AF), as detailed in Table 1-Left. The results reveal that general continual learning methods initially achieve satisfactory accuracy but suffer severe performance degradation on previous tasks as training progresses, resulting in higher forgetting rates. In contrast, continual forgery detection approaches exhibit more stable performance but still suffer from forgetting. It is worth noting that many of the compared methods Pan et al. (2023); Cheng et al. (2024) utilize a replay set, while our method strictly isolates all subtask data and still achieves superior anti-forgetting and state-of-the-art performance.
@@
Experiments on Manipulation Types Incremental Protocol. The dataset incremental protocol tested above demonstrates our method’s superior performance and anti-forgetting capability. Considering potential issues in the dataset incremental protocol, such as the domain gaps between different datasets and the possibility of shared manipulation methods across datasets, we perform experiments under the manipulation types incremental protocol [Hybrid, FR, FS, EFS], employing AUC as the primary metric for comparison. The experimental results are presented in Table 1-Right. The results show general continual learning approaches (LwF Li and Hoiem (2017), iCaRL Rebuffi et al. (2017a), DER Yan et al. (2021a)) exhibit unstable performance with high forgetting rates. Even specialized forgery detection methods demonstrate initial training instability – DFIL Pan et al. (2023) shows a 40.72% performance drop on previous tasks during FR-type training. Our method maintains stable learning throughout the entire training sequence, achieving both the highest final performance and the lowest forgetting rate. Comparative analysis reveals significant advantages: our approach surpasses SUR-LID Cheng et al. (2024) by 0.57% in average AUC, with 0.55% lower forgetting rate, and outperforms HDP Sun et al. (2024) by 16.64% AUC margin while reducing forgetting rate by 20.21%. These results validate our method’s performance against manipulation-type shifts in continual learning scenarios.
@@
Long-sequence continual learning experiment. To further validate the effectiveness of our proposed method on more manipulation types and extended task sequences, we selected 10 datasets from DF40 and completed long-sequence continual learning experiments under the Task10 protocol. We provide more details in the Appendix C.3. The average AUC metric and forgetting rate are plotted as shown in Fig. 3. The experimental results indicate that in long-sequence learning, both DFIL Pan et al. (2023) and HDP Sun et al. (2024) exhibit unstable catastrophic forgetting. At the same time, our method achieves the highest average AUC and the lowest forgetting rate with stable metrics.
@@ 4.3 Ablation Study
Effect of $\mathcal{L}_{\text{ort}}$ and $\mathcal{L}_{\text{llb}}$. We employ orthogonal loss $\mathcal{L}_{\text{ort}}$ to constrain both the subspace and gradient space of the currently learned LoRA and use $\mathcal{L}_{\text{llb}}$ to allocate the Real-LoRA and Fake-LoRAs. We conduct ablation studies on both losses, with the experimental results presented in Table 4. $\mathcal{L}_{\text{olora}}$ denotes the use of orthogonal subspaces without orthogonal gradients to validate the effectiveness of our orthogonal gradients; “All" represents the model’s performance incorporating all three losses. The results reveal that the $\mathcal{L}_{\text{ort}}$ significantly enhances the model’s anti-forgetting capability. Meanwhile, $\mathcal{L}_{\text{llb}}$ primarily improves the model’s Average Accuracy (AA) by facilitating task-specific allocations among LoRAs and maintaining an importance-weighted response matrix for collaboration.

### v3

@@ 4.1 Experimental Setup
Datasets. To facilitate comparisons, we select multiple datasets, employing different forgery methods to manipulate and generate faces from distinct domains. FaceForensics++ (FF++) ffpp contains 1K real face videos. These faces are manipulated by four forgery methods, resulting in a total of 4K fake videos. The Deepfake Detection (DFD) dfd dataset comprises over a hundred real face samples and more than 1K forged face samples. For the Deepfake Detection Challenge (DFDC) dfdc dataset, we utilized the Preview version (DFDC-P), which includes 5K videos and two forgery methods. Celeb-DF v2 (CDF2) celeb comprises 590 real and 5,639 fake videos. DF40 df40 is a large-scale dataset comprising 40 types of forgery methods and over 100K fake videos. We select [FF++, DFDC-P, DFD, CDF2] as our dataset-incremental testing protocol. From DF40, we choose three forgery methods corresponding to different forgery types: Face-Swapping (FS): BlendFace, Face-Reenactment (FR): MCNet, and Entire Face Synthesis (EFS): StyleGAN3. The datasets of selected forgery methods and a base dataset with hybrid forgery types FF++ jointly compose our forgery-type-incremental testing protocol: [Hybrid, FR, FS, EFS].
Implementation Details and Metrics. We utilize a pre-trained ViT model as the initial backbone. The pre-training data is sourced from real faces in the FF++ dataset. We conduct self-supervised pre-training through a data augmentation approach sbi. Regarding data utilization, we strictly isolate data across tasks, including real and fake faces. Similar to DFIL dfil, we randomly sample 100 videos from the current dataset and train for 20 epochs in each task, maintaining a balance between real and fake samples. The hyperparameter settings are adjusted according to the number of training epochs, and detailed information is provided in Appendix C.2. We employed the Adam optimizer with parameters set to $\beta_{1}=0.9,\beta_{2}=0.999$. The learning rate is set to 1e41e-4, and the batch size is set to 128. The rank (subspace dimension) for LoRA was determined using a grid optimization algorithm similar to InfLoRA inflora. We utilize Accuracy (Acc) and Area Under the Curve (AUC) to evaluate the overall performance of the model. Average Forgetting (AF) indicates the model’s ability to retain information from previously learned tasks. Let $a_{t,i}$ denote the evaluation $Score$ (Acc or AUC) on task $i$ after the model has been trained up to task $t$. The $AF_{T}$ (for $T>1$) is defined as:
$AF_{T}=\frac{1}{T-1}\sum_{i=1}^{T-1}\left(a_{i,i}-a_{T,i}\right),$ (15)
where $a_{T,i}$ is the score on task $i$ after learning the final task $T$.
@@ 4.2 Comparison Experiments
Experiments on Dataset Incremental Protocol. To comprehensively evaluate the performance of the proposed method, we conduct comparative experiments with general continual learning approaches (e.g., LwF lwf, ER er, and SI si) and specialized continual face forgery detection methods (e.g., CoReD cored, DFIL dfil, and DMP dmp). We ensure that the improvements stem from the new strategy and maintain the fairness of the experiments, as detailed in Appendix C.1. Under the dataset incremental protocol, the model is sequentially trained on task sequences constructed from [FF++, DFDC-P, DFD, CDF2]. After completing each task’s training phase, we evaluate model accuracy on both the current and previously learned tasks. Quantitative results are organized in a lower triangular accuracy matrix, accompanied by two key metrics: average accuracy (Avg) and average forgetting (AF), as detailed in Table 1-Left. The results reveal that general continual learning methods initially achieve satisfactory accuracy but suffer severe performance degradation on previous tasks as training progresses, resulting in higher forgetting rates. In contrast, continual forgery detection approaches exhibit more stable performance but still suffer from forgetting. It is worth noting that many of the compared methods dfil; surlid utilize a replay set, while our method strictly isolates all subtask data and still achieves superior anti-forgetting and state-of-the-art performance.
@@
Experiments on Manipulation Types Incremental Protocol. The dataset incremental protocol tested above demonstrates our method’s superior performance and anti-forgetting capability. Considering potential issues in the dataset incremental protocol, such as the domain gaps between different datasets and the possibility of shared manipulation methods across datasets, we perform experiments under the manipulation types incremental protocol [Hybrid, FR, FS, EFS], employing AUC as the primary metric for comparison. The experimental results are presented in Table 1-Right. The results show general continual learning approaches (LwF lwf, iCaRL icarl, DER der) exhibit unstable performance with high forgetting rates. Even specialized forgery detection methods demonstrate initial training instability – DFIL dfil shows a 40.72% performance drop on previous tasks during FR-type training. Our method maintains stable learning throughout the entire training sequence, achieving both the highest final performance and the lowest forgetting rate. Comparative analysis reveals significant advantages: our approach surpasses SUR-LID surlid by 0.57% in average AUC, with 0.55% lower forgetting rate, and outperforms HDP hdp by 16.64% AUC margin while reducing forgetting rate by 20.21%. These results validate our method’s performance against manipulation-type shifts in continual learning scenarios.
4.3 Long-sequence Continual Learning Experiments
@@
To further validate the effectiveness of our proposed method on more manipulation types and extended task sequences, we selected 10 datasets from DF40 and completed long-sequence continual learning experiments under the Task10 protocol. We provide more details in the Appendix C.3. The average AUC metric and forgetting rate are plotted as shown in Fig. 3. The experimental results indicate that in long-sequence learning, both DFIL dfil and HDP hdp exhibit unstable catastrophic forgetting. At the same time, our method achieves the highest average AUC and the lowest forgetting rates.
4.4 Motivation Validation Experiments
Figure 4: Motivation Validation. We track the variation range of the orthogonal gradient loss during the training process for two orthogonal sequences, one modeling real faces and the other modeling forged faces. For each sequence, the two lines represent the upper and lower bounds of the loss, and the shaded area between them indicates its range of variation. We observe that real faces exhibit an obvious smaller orthogonal gradient loss.
Our proposed method is based on a key observation: abundant and unbiased real faces exhibit a more compact distribution across datasets than forged faces created by various methods. Building upon a t-SNE experiment that validates this observation, we further confirm that real faces occupy a more similar input space than forged faces within our proposed framework. We first theoretically demonstrate that a more similar input space corresponds to a smaller orthogonal gradient loss. Specifically, when a new orthogonal LoRA is configured to model a highly similar domain, this domain will possess a highly similar input space $\mathbf{H}_{t}\in\mathbb{R}^{d_{I}\times n}$. From Eqn. 7, it can be deduced that the resulting $\mathbf{V}^{T}_{t}$ will also be highly similar. Consequently, in Eqn. 8 and Eqn. 9, the term $\sum\left\|\mathbf{G}_{i,t}\right\|^{2}$ yields a near-zero value because its components are approximately orthogonal.
Based on this, we design a new experiment consisting of two separate orthogonal LoRA sequences: one for modeling real faces and the other for modeling forged faces. These sequences learn from the task series [FF++, DFDCP, DFD, CDF2]. We record the range of orthogonal gradient loss values for the two sequences during the training process to investigate the cross-dataset similarity of their respective input spaces. The experimental results are presented in Fig. 4. Starting from the second task, the orthogonal gradient induced by the real face data for a new LoRA is smaller than that induced by the forged face data. The results suggest that real faces from different datasets share a similar input space $\mathbf{H}_{t}$, leading to an orthogonal gradient loss that is an order of magnitude smaller. Moreover, the orthogonal LoRA sequence introduces additional computational overhead compared to a single Real-LoRA. This experiment mutually validates our initial t-SNE results, demonstrating that using a single, shared Real-LoRA to model real faces is both reasonable and necessary.
@@ 4.5 Ablation Study
Effect of $\mathcal{L}_{\text{ort}}$ and $\mathcal{L}_{\text{llb}}$. We employ orthogonal loss $\mathcal{L}_{\text{ort}}$ to constrain both the subspace and gradient space of the currently learned LoRA and use $\mathcal{L}_{\text{llb}}$ to allocate the Real-LoRA and Fake-LoRAs. We conduct ablation studies on both losses, with the experimental results presented in Table 4. $\mathcal{L}_{\text{olora}}$ denotes the use of orthogonal subspaces without orthogonal gradients to validate the effectiveness of our orthogonal gradients; “All" represents the model’s performance incorporating all three losses. The results reveal that the $\mathcal{L}_{\text{ort}}$ significantly enhance the model’s anti-forgetting capability. Meanwhile, $\mathcal{L}_{\text{llb}}$ primarily improves the model’s Average Accuracy (AA) by facilitating task-specific allocations among LoRAs and maintaining a weighted response matrix for collaboration.
@@ 5 Conclusion
Acknowledgement
This work was supported in part by Chinese National Natural Science Foundation Projects U23B2054, 62276254, 62176256, Beijing Natural Science Foundation L242092, the Science and Technology Development Fund of Macau Project 0140/2024/AGJ.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state the proposed claims and motivations in the abstract and introduction, and list our contributions at the end of the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a detailed discussion of the limitations of this work in Appendix Section: Limitations, including the potential increase in model scale with training.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section: Methodology of the paper, we provide the full set of assumptions and a complete (and correct) proof for each theoretical result, such as the reasons why orthogonality constraints cause forgetting, the derivation of orthogonal gradients, and the derivation of the label-guided localized balancing strategy.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In the Experimental Setup section of Section: Experiments, we present all experimental details for reproducibility, including the datasets used, the backbone network, training details, hyperparameters, and the unique task settings and task sequences specific to continual learning. In Section: Methodology, we provide a detailed description of all aspects of our method and the loss functions. Furthermore, we will make all training code and training details publicly available.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will release all our training code, inference code, experimental settings, and trained models for public availability. The datasets used in the paper may require additional application and access, such as FaceForensics++. We will provide the official link to these datasets, as well as our preprocessing code for handling them, to ensure full reproducibility of the results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In the Experimental Setup section of Section: Experiments, we present all experimental details for reproducibility, including the datasets used, the backbone network, training details, hyperparameters, and the unique task settings and task sequences specific to continual learning.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The experiments are based on two metrics: accuracy and AUC. First, due to our limited experimental resources, conducting repetitive experiments is impractical. Second, in the field of Face Forgery Detection, error bars or statistical significance tests are not required. None of the compared Face Forgery Detection methods provide statistical significance tests or error bars in their papers.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In the Experimental Setup section of Section: Experiments, we provide the computational resources required for the experiments, and in the appendix, we present information related to computational resource consumption, including the total number of model parameters, the number of backbone parameters, and the trainable parameter count.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: In the paper, we carefully adhere to the NeurIPS Code of Ethics in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Undoubtedly, the work presented in this paper has positive societal impacts, such as preventing fraud and identifying fake information, which we discuss in detail in the introduction.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release any pretrained language models, image generators, or scraped datasets that could potentially be misused, thereby ensuring that there are no such risks associated with the paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have provided citations and application conditions for the datasets and methods used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We do not propose any new dataset-related assets, and the trained models as well as all code will be open-sourced.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: There are no crowdsourcing experiments or research involving human subjects in the paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: There are no crowdsourcing experiments or research involving human subjects in the paper.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method in the paper does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.